Volume 1, Number 3 October 1992

TEX and TUG NEWS

for and by the TEX community

A Publication of the TEX Users Group Electronic version TEX and TUG NEWS

Mission Statement

The TEX Users Group (TUG) provides leadership:

1. to encourage and expand the use of TEX, , and related systems

2. to ensure the integrity and portability of TEX, , and related systems

3. to foster innovation in high-quality electronic document preparation

TEX and TUG NEWS is a newsletter for TEX and L^{A}TEX users alike: a forum for exchanging information, tips and suggestions; a regular means of communicating news items to one another; a place where information about TEX and TUG can be quickly disseminated.

Throughout the newsletter “TEX” is understood to mean TEX, L^{A}TEX, AMS-TEX, and other related programs and macros. TEX and TUG NEWS is produced with the standard L^{A}TEX distribution, and is to be as portable a document as possible.

The entire contents of this newsletter are being placed in the public domain. The source file of this issue will be placed in the aston, shsu, and stuttgart archives, as well as at the heidelberg, labrea, and ymir archives. Copying and reprinting are encouraged; however, an acknowledgement specifying TEX and TUG NEWS as the source would be appreciated.

Submissions to TEX and TUG NEWS should be short, the macros must work, and the files must run without special font or graphics requirements: this is to be a portable newsletter (the new font selection scheme has not yet been implemented). Correspondence may be sent via e-mail to tug@math.ams.org with the subject line NEWSLETTER. Regular mail should be addressed to the Publications Committee, c/o TUG Office at the address below.

TEX Users Group  Phone: 401-751-7760
658 North Main Street  FAX: 401-751-1071
P.O. Box 9506  E-mail: tug@math.ams.org

Providence, RI 02940-9506 USA

TEX is a trademark of the American Mathematical Society

TEX and TUG NEWS (ISSN 1065-240X) is published four times a year, February, July, October and December, by the TEX Users Group, 653 North Main Street, P.O. Box 9506, Providence, RI 02940, U.S.A. The 1992 dues for individual members are $60.00 for regular members and $50.00 for students.

Second-class postage paid at Providence, RI and additional mailing offices. Postmaster: Send address changes to the TEX Users Group, P.O. Box 9506, Providence, RI 02940, U.S.A.

Membership in the TEX Users Group is for the calendar year, and includes all issues of TUGboat and TEX and TUG NEWS for the year in which membership begins or is renewed. Individual membership is open only to named individuals, and carries with it such rights and responsibilities as voting in the annual election.

Editorial

Well, here we are with issue no. 3 of the year, and one more to go. Despite a rather slow start to the year, TEX and TUG NEWS will indeed be a quarterly publication; we are attempting to achieve a staggered schedule of publication in 1993 with respect to TUGboat, with issues appearing in January, April, July (just before the Annual Meeting), and October. For this year, we’re hoping to have the last issue out for December.

This means we have room for lots of material from you! One important area to my mind which ought to be explored is how those of you who have to provide support to users do that. Often it seems just a case of catch-as-catch-can, but is there a better way? What about binders with examples (source + output)? What about local archives? Templates of common document styles?^{1} Mimi Burbank, editor of the 1992 TUG Proceedings, has some ideas she’s implemented at SCRI (p. 14). Why not send in your own TEX tactics?

Speaking of Proceedings, the Annual Meeting just held in Portland, Oregon was a great success, in my estimation (more on p. 19). And we also had the first TUG fundraiser — a bowling tournament! Find out more on p. 21. The Proceedings have just been mailed out, and should be in your hands soon.

And big news about next year’s meeting: TUG’93 will be in England, at Aston University, in the city of Birmingham. Mark July 26–29 on your 1993 calendars for the “World Wide Window on TEX”.

In the last editorial, I said that the winner of the Knuth Scholarship would be announced at Portland — and she was. Congratulations, Jenny Smith! Information on applying for the 1993 Knuth Scholarship is provided in this issue (p. 3). As well, there’s more information on the upcoming election of a new TUG President (p. 25).

Membership renewal notices are being sent out soon, so be on the lookout for them, especially if you receive your TUGboats via an institutional membership. Make sure they renew for you! Along with the renewal is an invitation to provide some additional support for the L^{A}TEX3 project — on p. 4 we have an open letter from the L^{A}TEX3 team to all TUG members and TEX users. As well, they are now looking for L^{A}TEX users to lend a hand in carrying out additional and necessary research to aid in their efforts. See p. 18 for preliminary information, with more to come in TUGboat 13, no. 4.

The L^{A}TEX3 project is actually just one of several which come under the umbrella of TUG’s Technical Council — Michael Ferguson has written this issue’s feature article on the various working groups currently doing research (pp. 5–8). Go have a look and see what’s happening — it’s pretty interesting.

1Templates were in fact the subject of a paper by Hope Hamilton at the 1989 TUG meeting at Stanford. See TUGboat 10, no. 4, pp. 541–548. And the title of a paper by Anita Hoover seems to deal with the subject in general — see the Prague Proceedings contents, p. 28.

This issue also carries a list of all TUG courses scheduled so far for 1993 (p. 26). A recent TEX for Publishers course (September 30 in New York) was a resounding success with 28 attendees, and more hopefuls who had to be turned away.

And some notes on changes:

1. TUG’s e-mail address has been slightly modified now so that .com has been

     replaced by .org: tug@math.ams.org.

2. The New Publications column will now be handled by Peter Schmitt, Institute of Mathematics, University of Vienna. If you hear of any new books on TEX, L^{A}TEX, or , send him a note (a8131dal@helios.edvz .univie.ac.at); and don’t forget that we also list the latest issues of newsletters from other TEX user groups. As for reviews of books already cited here, TUGboat 13, no. 2 has reviews of Beccari (listed in TTN 1,1), and reviews of the Hahn and Eijkhout books (listed in TTN 1,2). If you want to review any books we list, send a query to tugboat@math.ams.org.

3. Jackie Damrau (Superconducting Supercollider Lab., Texas) and Chris Carruthers (University of Ottawa) are now lending a hand in getting TTN out to you. Chris will also be handling the electronic distribution of TTN’s files to the various archives which store TEX material.

Another change was sparked by a suggestion from Karl Berry: could we please set up all articles so that authors and their affiliation are part of the titleblock area, rather than having them at the top of some articles and at the end of others. So we’ll try that one out — it does make it easier on the reader to get all the information in one place. … although I don’t think I’ll change policy on the editorial page, Karl!

One final note here: I’m looking for some humour! TTN could use some light-hearted content occasionally — Peter Flynn’s made a start in his column, but I want more! If you have ideas, or if you’d like to look after this occasional TTN item, drop me a line.

Christina Thiele

Editor, TEX and TUG NEWS

Did you know …?

… that as of October 1st, TUG membership stood at 3,172. And slightly less than half of that is from outside the United States (which counts 1,719 members): 1,053 TUG members are in Europe, 171 in Canada and Mexico, and 229 elsewhere around the globe.

The Donald E. Knuth Scholarship

The 1992 Donald E. Knuth Scholarship has been awarded to Jenny Smith, an employee of the Production Department of John Wiley and Sons, Ltd., Chichester, England. Ms. Smith was honored at the 1992 TUG Annual Meeting in Portland, Oregon; she is the seventh recipient of the Scholarship.

The intent of the Knuth Scholarship is to encourage the increase of knowledge about TEX and to sharpen the TEX skills of non-technical users.

The 1993 competition will be open to all 1993 TUG members holding support positions with duties that are secretarial, clerical or editorial in nature. It is therefore not intended for those with a substantial training in technical, scientific or mathematical subjects and, in particular, it is not open to anyone holding, or studying for, a degree with a major or concentration in these areas.

The award will consist of an expense-paid trip to the 1993 TUG Annual Meeting at Aston University in Birmingham, England, and to the Scholar’s choice from the short courses offered in conjunction with that meeting. A cap of $2,000 has been set; however, this does not include registration fees for the meeting and short course, which will be waived.

Schedule

March 1, 1993  Deadline for receipt of submissions
March 15–May 10, 1993 Judging period
May 17, 1993  Notification of winner

An announcement in TUGboat 13, no. 3, pp. 395–396 contains a description of the material that must be submitted by candidates, which includes a TEX project of the candidate’s choice, to be judged by the Scholarship Committee, and certain personal information.

The 1993 Scholarship Committee consists of

Chris Rowley, Open University, U.K. (Chair)

David Salomon, California State University, Northridge Jenny Smith, John Wiley and Sons, Ltd., Chichester, England

Copies of the TUGboat article and some additional information are available from the TUG Office. To obtain a copy, write to TUG (address on inside front cover of this issue), or send a message by e-mail to TUG@Math.AMS.org with the

subject Knuth Scholarship request.

Barbara Beeton, Outgoing Liaison Nico Poppelier, New Liaison

L^{A}TEX3 Update

To Members of the TEX and TUG Community:

Shortly, current members of TUG will receive their membership renewal forms. On that form, you will notice a line requesting your (optional) contribution to the L^{A}TEX3 project. We hope this letter will explain that item to you.

The L^{A}TEX3 Project is an attempt by a small group of volunteers to revise the current version of L^{A}TEX in the light of almost eight years of intensive experience on the part of the user community. This is not the place to write about this project.^{1} Readers may be assured, though, that we are committed to retaining the ease of use and the logical document design of the “old” L^{A}TEX, while incorporating features that people have most missed.

It’s appropriate to mention two other aspects of the new L^{A}TEX. L^{A}TEX3 will remain in the public domain, and will be fully documented at all levels.

Public domain software is not free software, although no money goes toward salaries — we all work as volunteers. The development team is fully international, and although we can accomplish many things via email, there are times when face to face meetings must take place. Furthermore, certain testing must take place using equipment that none of us may currently have access to.

It is for these reasons that we ask for financial support from you. Any sum will be gratefully appreciated. The amount need not be large. Contributions of state of the art equipment are also needed.

Thanks in advance for your generosity. We look forward to hearing from you.

Frank Mittelbach Chris Rowley

for the L^{A}TEX3 Development Team Alan Hoenig

L^{A}TEX3 Liaison for TUG

More L^{A}TEX3 News!

Turn to p. 18 for details on how you can volunteer to lend a hand with the L^{A}TEX3 project; and on p. 21, there’s more: a fundraising event held at the TUG meeting in Portland this past July.

1For additional information, see, for example, articles which have appeared in TTN 1,1 and 1,2 of this year, and in TUGboat 13, no. 1, pp. 96–101. As well, information presented at TUG’92 in Portland this past July will be appearing in the Proceedings issue of TUGboat.

The Technical Council

Michael J. Ferguson Chair, Technical Council

                                      mike@inrs-telecom.uquebec.ca

The TUG Technical Council was established, in its present form, at the February 1992 meeting of the TUG Board of Directors at Aston University in Birmingham, England, and is expected to serve as the focal point for technical activities within TUG. The Technical Council does not intend to actually carry out the technical activities but to act rather as a facilitator for these activities. The actual technical work will be carried out by Technical Working Groups (TWG), each with its own chair, who will have complete control over the technical activities of the group. A member of the Technical Council (TC) will act as a liaison with the TUG Board and hence with the TUG membership.

The Technical Council Members are as follows:

        Michael Ferguson (Chair): mike@inrs-telecom.uquebec.ca
        Alan Hoenig: ajhjj@cunyvm.cuny.edu
        Yannis Haralambous: yannis@gat.citilille.fr

Donald Knuth — Grand Wizard (Honorary)

Hermann Zapf — Grand Wizard of Fonts (Honorary)

Technical Working Group Structure

The Technical Working Group (TWG) is most important. It does all the technical work and normally consists of a TWG Chair, TWG Members, and a Technical Council Liaison. Some guidelines for a TWG are

• The Technical Working Group Chair is appointed by the Technical Council. The Working Group Chair is solely responsible for the study procedure and structure of the Working Group. All Technical Working Group members serve at the discretion of the Working Group Chair.

• The Technical Council Liaison is responsible for timely reporting of Working Group activities to the Technical Council, and for any Working Group Reports or Recommendations that the TWG Chair would like to be officially issued by TUG.

• The Technical Council will determine the mandates of Technical Working Groups.

• Each Technical Working Group will make a report on its activities at each meeting of TUG’s Board of Directors via the Technical Council and the TWG’s Liaison.

In the few months that the Technical Council has been in existence, we have identified a need for several types of TWGs. These are as follows:

Independent Research Projects (IRP-TWG):

There are several ongoing research projects of vital interest to the entire TEX community. The IRP-TWG is designed to be a home for these projects. Currently there are two projects associated with WG-92-00: Independent Projects: L^{A}TEX3 and NTS — New Typesetting System.

Standards and Coordination (SC-TWG):

The primary intent of a SC-TWG is to evaluate technical issues in a specific area and to recommend actions so that a stable, consistent, and effective environment for document preparation and production may be maintained. Some examples of this type are WG-92-01^{1} on TEX Extended Mathematic Font Encoding and WG-92-05 on TEX Archive Guidelines.

Special Interest (SI-TWG):

A Special Interest TWG is intended to serve as a focus for those individuals that are interested in issues related to applying TEX in a specific areas. These TWGs are intended to be made up of ordinary TEX users and could become quite large. At present WG-92-04 TEX for the Disabled is our only SI-TWG but we encourage other individuals that would like to organize a new SI-TWG to contact any member of the Technical Council.

Technical Working Groups

WG-92-00 (IRP-TWG) Title: Independent Projects

Mandate: To recognize and report to the TUG Board and the TEX Community on important projects which are independent of TUG but are of concern to the entire TEX Community. Due to its independent nature, each such project will retain complete control over its membership, working methods and agenda. The chair of the Independent Projects Working Group shall report, with the cooperation of a project’s management, on that project’s activities to the TUG Board and thence to the TEX Community. The reporting function of the TWG Chair shall in no way interfere or preclude reporting or dissemination of the results of a project by the project management or by members of the project.

TWG Chair: Alan Hoenig WG-92-01 (SC-TWG) Title: TEX Extended Mathematic Font Encoding

1The numbers and titles are explained in the next section.

Mandate: To create font encoding standards for Mathematical fonts used in TEX systems.

TWG Chair: Barbara Beeton, bnb@math.ams.com

TC Liaison: Michael Ferguson

WG-92-02 (SC-TWG) Title: System Implementor Coordination

Mandate: To facilitate the organization, reporting, and dissemination of ports of TEX and its related systems on various computing platforms. The Technical Working Group shall consist of a “System Coordination Chair” for each different computing platform. Each coordinator shall be responsible for the organization and dissemination of information related to the porting of TEX and its systems to their platform. Preparation of the software for distribution shall be the responsibility of the System Implementor.

TWG Chair: Michael Ferguson

Current System Coordination Chairs (August 1992):

• IBM VM/CMS: Joachim Lammarsch,

x92@vm.urz.uni-heidelberg.de

• VAX/VMS: David Kellerman, kellerman@nls.com

• UNIX: Karl Berry, karl@cs.umb.edu

WG-92-03 (SC-TWG) Title: Multiple Language Coordination

Mandate: The primary purpose of this working group is to obtain, for TEX systems, a consistent means for implementing, accessing, and describing, the fonts, ligature rules, hyphenation patterns and other special requirements for a given linguistic group. Each linguistic group shall have a Group Coordination Chair who shall be responsible for discussions and implementation of those features and macros required for their group. In addition, the working group shall strive to create uniform mechanisms for access and definitions that are common to each linguistic group.

TWG Chair: Yannis Haralambous

Common Language Issues: Johannes Braams, IL-Braams@pttnrl.nl Current Group Coordination Chairs:

• African: Jörg Knappen, knappen@vkpmzd.kph.uni-mainz.edu
• Greek: Silvio Levy, levy@math.berkeley.edu

• Scandinavian: Jan Michael Rynning, jmr@nada.kth.se

WG-92-04 (SI-TWG) Title: TEX for the Disabled

Mandate: The primary purpose of this working group is to serve as a forum for those people interested in using and/or enhancing TEX to serve the needs of the disabled.

TWG Chair: T.V. Raman, raman@cs.cornell.edu

TC Liaison: Michael Ferguson

WG-92-05 (SC-TWG) Title: TEX Archive Guidelines

Mandate: The purpose of this Technical Working Group is to develop guidelines for the effective management and utilization of major TEX archives, and to initiate communication among the maintainers of the existing archives for the purpose of coordination and synchronization. Areas to be considered include, but are not limited to:

• what physical resources are needed

• what kind of management support is desirable

• what personnel resources are necessary and how should they be organized

• breadth and depth of coverage, general and specific subject archives

• strategies for organizing and managing the archive holdings

• strategies for archive synchronization

• dissemination strategies of archive contents

• should there be “official” archives

TWG Chair: George Greenwade, bed gdg@SHSU.edu

TC Liaison: Alan Hoenig

New Publications^{†}

Jiří Zlatuška, ed. EuroTEX 92, Proceedings of the 7th European TEX Conference. Brno: Masaryk University and Prague: Czechoslovak TEX Users Group (CsTUG), 1992. ISBN 80-210-0480-0.

• Copies can be ordered directly from CsTUG for the price of DM 30.– plus postage (or the equivalent in some other currency) at the following address: Československé sdružení uZivatelů TEXu (CsTUG), Matematický ústav UK, Sokolovská 83, CS-186 00 Praha 8, Czechoslovakia. They may also be available through your local TEX users group.^{1}

Stephan von Bechtolsheim. TEX in Practice. New York: Springer Verlag, Oct. 1992. 4-volume set (hardcover), US$169.00. ISBN 0-387-97296-X. Volumes also available individually, US$49.00 each:

Vol. 1: Basics. Pp. 359, 9 illus. ISBN 0-387-97595-0.

Vol. 2: Paragraphs, Math, and Fonts. Pp. 384. ISBN 0-387-97596-9. Vol. 3: Tokens, Macros. Pp. 544, 22 illus. ISBN 0-387-97597-7. Vol. 4: Output Routines, Tables. Pp. 300. ISBN 0-387-97598-5.

• “A recent surge of good TEX implementations for PCs has put TEX on the disks of many people including writers, designers, desktop publishers, and engineers. With such increased interest in TEX, there is a need for good TEX books. TEX in Practice is the ideal reference and guide for the TEX community. The four-volume set is written by an acknowledged expert in the field and addresses the needs of the TEX novice to the more experienced ‘TEXpert.’ The book provides step-by-step introduction to the various functions of TEX with many relevant examples.”

Stefan Schwartz and Rudolf Potuček. LATEX — Satzkunst statt DTP. [“L^{A}TEX — The Art of Typography rather than DTP”]. (Special issue in the series CHIP-Anwender-Praxis.) 1992. DM 34,– . ISBN 3-8023-1178-7.

• Can be obtained from: Vogel Verlag, Postfach 67 40, D-8700 Würzburg, Germany.

Latest issues of other TEX newsletters

Les Cahiers GUTenberg: no. 13 juin 1992, 90pp. ISSN 1140-9304.

†The aim of this column is to provide up-to-the-minute information on all publications on TEX, , and related topics. Descriptions are usually taken from promotional materials, and should therefore not be considered an evaluation of the book. Reviews may appear later in TUGboat.

1The Table of Contents of the Proceedings appears on p. 28 of this issue, and a short report on the conference by Alan Hoenig appears on pp. 22–23.

Typographer’s Inn

Peter Flynn

University College Cork

                                          cbts8001@iruccvax.ucc.ie

At the annual meeting in Portland (reported elsewhere in this issue) there was the chance to discuss typography with old and new friends. Some of it centered around the differences between European and North American typographic practice, and raised more than a few hackles.

And look! you get sectioning in this column, too ◯ ⊙⊙ ; △ ; ⌣

Punctuation and quotes

I recall being taught (at six or seven years of age) that punctuation after quotation was always enclosed within the closing quotes, “like this.” This was in England, and I always felt it was wrong, because the fullpoint was not an integral part of the quoted text, but a part of the surrounding sentence. The USA (and Canada?) still adheres to this mode, but recent practice this side of the puddle is now to put the punctuation after the closing quotes. Nevertheless, I still feel the sense and usage of the mark should determine its position, not an arbitrary rule. Does anyone have burning feelings on this?

Bits & pieces

My mailbox has been stuffed with criticism and good advice. I’ve tried to answer them all, but about a dozen were trashed in a crash, I’m afraid. Thomas Reid (Texas A&M) picked up on my point that a common gripe about TEX is “it only has one typeface”, and asked how many other typefaces available to TEX have matching small caps fonts? Of the ones I use, only the MetaFoundry version of Times has this, but doubtless there are more, especially among all the Monotype fonts now becoming available in outline form (an old favorite, Mono’s Oldstyle Prumyslava, will alas probably never appear in outline). Tom also mentioned using Optima fonts, which do not have matching small caps. Sans-serif faces don’t, and although Optima is probably best described as semi-serif, I have never felt the lack of small caps in it, because of the large x-height.

Kathleen Lyle from Sheffield says her authors complained when a small-caps word was pluralised with a lowercase ‘s’. I actually do this habitually and like it: has anyone else had this problem? She also takes me to task for (wrongly?) using an en-rule in “ftp–accessible”. My reason was the change of font from \tt to \rm but this is probably a personal quirk of mine.

Underlining

Victor Eijkhout brought up the subject of underlining, which (we are all taught) is very naughty in decent typography, being regarded as a hangover from the days of typewriters. He cites Updike’s Printing Types and Tschichold’s Die Neue Typographie as using underlined headings, and I partly agree: an experienced typographer can probably get away with all sorts of tricks you or I would probably not attempt. I think in many cases, the ‘underlining’ is often a spaced rule, a typographical entity in its own right, despite what GBS said (TEXbook, p. 229), and not strictly underlining, which in hot metal was impossible without specially-cast sorts.

Antiquarianism?

Several people downloaded my version of the WWP’s The Ladies’ Defence and have asked how close it was to the original page layout and original fonts. The answer is, not at all: I have never even seen the original. I wanted a modern edition, but retaining some of the flavour of 17th century printing. Hence I have obeyed in all detail the encoding which the WWP inserted for the use of italics and boldface, exactly as it was originally printed. Probably not a very clever thing to use CMR for it, but infinitely better than any of the ‘35’ (the Adobe PostScript suite).

Page layout

A matter of page layout was brought to me the other day: a local user, whose work involves mainly pamphlets and brochures (currently done with a wordprocessor), habitually designs each double-page spread individually and separately, and wanted to know how TEX might help. My feeling was that if there was only a small selection of variant layouts, they could be encoded into macros, but this turned out not to be so. Reluctantly it seemed to be a case where a visuallyinteractive system, which allows on-screen manipulation of the grid and type, is better suited to the task. Of course TEX can do the job, but for a user with good design appreciation but little time to spare to learn what would be quite a lot of TEX internals, I felt one of the more advanced wordprocessors would be more suited. I realise that to a lot of TEXies this is a heresy of the first order, but what would have been your recommendation?

As we go to press, a reminder on the network from Yannis Haralambous that a program was shown at the Paris conference last year which lets you interactively edit a .dvi file, inserting graphics, changing the position and order of things. It’s in β–test at the moment, but details can be got from Maurice Laugier at louijean@cicg990.grenet.fr—as Yannis said to me later, maybe this is one more chance for TEX to survive in a wysiwyg world!

“Hey — it works!”

Marking lines in the margin

Peter Schmitt University of Vienna

                               a8131dal@awiuni11.EDVZ.UniVie.ac.at

Putting some material into the margin is a standard procedure. Usually each margin note is handled separately, but sometimes this may be inconvenient, as proven by the following question which was recently raised on the German TEX list:

Is there a convenient way to mark a number of lines by asterisks placed into the margin?

One possible solution is provided by the definitions of \marginmarkson and \marginmarksoff, shown below (of course, shorter control sequence names could be used!):

   \long\def\readword#1 {\strut\putmark#1 \next\relax}
   \def\margmark{$\ast\hskip10pt$}
   \def\putmark{%
       \vadjust{\vbox to 0pt{\vss\noindent
                             \llap{\strut\margmark}\vskip0pt}}%
       }
   \def\marginmarkson{\global\let\next=\readword\next}
   \def\marginmarksoff{\global\let\next=\relax}

The basic idea is to read the text word by word (or, more precisely, from space to space) and — with each step — to use \vadjust to put an asterisk

∗ into the margin. This feature is turned on by \marginmarkson and turned off ∗ ∗ ∗ ∗ ∗ ∗

∗ by \marginmarksoff. The command \readword reads the material to the next ∗ ∗

space, sets the \vadjust (using \putmark), reinserts the material (including the space) and finally iterates the process. Since \vadjust is inserted for the first time with the word following the \marginmarkson, the first asterisk is put into the margin opposite the line that contains this word. The last line marked will be the line that contains the last word before \marginmarksoff.

These macros can be used independently of groups formed by \bgroup or \begingroup and allow (short) groups between braces (long groups may cause a missing margin mark).

In the definition of \readword, the \strut serves two purposes: it allows the use of \marginmarkson at the beginning of a paragraph since it guarantees that TEX will be in horizontal mode when the first \vadjust is inserted. And it protects against the (unlikely) case where a line does not contain letters which reach below the baseline since otherwise — in this case — the marginmark would be positioned higher than usual.

The \relax prevents braces enclosing some group from disappearing when \readword reads its argument.

Remarks:

• Clearly, the asterisk used is just an example. It can easily be replaced by giving \margmark another meaning, as in the following examples which yield various types of changebars:

\def\rule{\vrule\strut\hskip5pt}
\def\bar{\vrule height8.5pt depth 4.5pt width1.pt\hskip10pt}
\def\dbar{\vrule height8.5pt depth 4.5pt width1.pt\hskip5pt%
          \vrule height8.5pt depth 4.5pt width1.pt\hskip10pt}

The mark can be changed by \let\margmark=\bar even in the middle of a marked paragraph, provided one accepts that the line where the change takes places shows both marks (as shown to the left). If applicable, the definition of \margmark should use explicit fonts. Otherwise a font change in the text may influence the appearance of the mark.

• This definition might fail in situations using complicated constructions that interfere with \vadjust, but it works with ordinary text, both in TEX and L^{A}TEX. Another source of problems may be the use of catcode changes in marked text. One has to take care that the control sequence executing the change is read by \putmark before the argument; that is, one has to write

                read by\verb| \putmark| before

instead of

                read by \verb|\putmark| before

to obtain the phrase used above [see Editor’s Note, next page].

• Of course, the algorithm used is not very efficient, since it puts one mark into the margin for every word (space) in the line. But this does not matter since these marks are all put into exactly the same location. And it works since usually each line will contain at least one space.

Editor’s Note:

This problem came to light as I was preparing Peter’s article for TTN. I use L^{A}TEX’s \verb construction with vertical bars as delimiters, and wanted to demonstrate the macros in the actual text. This, however, resulted in some rather severe interference: the text in the in-line verbatim construction disappeared! Peter wrote: “This first puzzled me — but the reason is quite clear. It is due to the (well-known) fact that taking an argument fixes the catcodes. Since \putmark reads a word (up to and including the space right after it), it takes up the \verb|\let\margmark=\bar| (for instance) and the control sequences are stored away before \verb takes action: the assignment gets executed instead of printed.”

As well, he warns: “There is another problem inherent in the method used here: if the lineskip is not constant, then the margin marks will be distributed irregularly or, in the case of bars, there may be holes (e.g., with math). But, in my opinion, there is a place for quick and not completely general macros … As far as I know a general solution would require more complicated code.”

We thought this additional note on some problems related to the marginmarks macro would be instructive and useful: items offered in this column do not pretend to be the definitive solutions to situations which arise — I’ve said so from the beginning. They’re quick solutions, and in the general case, do work. But as this example shows, interference is always a possibility — and not just with hacks. Users should be prepared to wrestle occasionally with someone else’s code — this doesn’t invalidate the code, nor does it mean that borrowing macros should be discouraged. It’s just that one can’t be totally blind, or jump up and down when things don’t work. Be happy when they do — be philosophical when they don’t. Life’s too short to get exercised about the occasional problem.

When there’s only one of you and a hundred of them

Mimi Burbank

Supercomputer Computations Research Institute, Florida (SCRI)

Having been hired to provide TEX support for a group of 10–15 people, and having seen my job expand to providing support to hundreds of people located on many different systems, not to mention in many countries of the world, I have a great deal of sympathy for those of you who are new to this situation. If only someone had made suggestions on how to to simplify my labors. The following are some suggestions; basically, they represent my solutions to myriad variations on questions such as: “How do I do this?” or “Can you show me how to do this?” or “Where can I find out how to do this?”

• If you have files which are useful to many people, I recommend having your system manager set up a “public” directory — one which can be accessed by other users. In this directory, group similar files in subdirectories by topical area.

As an example, our group has a public disk which is divided into archives, conferences, database, manuals, proposals, preprints, and names of scientists (to name a few). Each directory is subjectively named and sub-divided appropriately. At present we have a total of 259 directories, and 4,464 files.

• If you have many people trying to access locally developed macros, consider adding their physical location to the search path of tex_inputs. Then a lot of questions can be conveniently answered with: “You can copy the

macro located at tex_inputs:file.name.”

Our proposals, annual reports, and vitæ generally use a standard set of macros for production, and we have added these to the tex_inputs so that scientists writing individual portions of the proposal can run their files in their home directories, rather than wait for me to run the files for them.

• After having a hundred different people ask the same question, and/or having several different scientists ask me to create the same bizarre symbol, I finally put these things together in one file which includes a description of the symbol, the macro defining it and the command calling the macro. This file is not to be \input; rather, it is to be copied and individuals may TEX the file themselves and either use, or “clone” a macro for their own personal use. If I have a new or odd macro, I place it in my file, annotated properly, and then tell everyone where to find the file.

Some TTN L^{A}TEX tricks

Christina Thiele

Once our macros settle down, there’ll be a write-up in TTN. But here are a few features specific to this issue which might be of interest, especially to those who are perhaps more familiar with TEX and not so comfortable with L^{A}TEX.

On page 28 of this issue, we have a common enough structure which isn’t all that easy to do in plain TEX, but which can be handled in a couple of ways with regular L^{A}TEX commands. The problem is to mix some text which spans the full width of a page with some text done in two columns.

L^{A}TEX’s \twocolumn command can be used if the bulk of a document is single-column (if the entire document is to be in two columns, then it is more appropriate to use the [twocolumn] option with the opening \documentstyle command).

\twocolumn[\Section{Euro\TeX~'92: Contents from the Proceedings}
\vspace{-6pt}
The following is excerpted from the {\sl Proceedings of the 7th
European \TeX\ Conference\/}, held in Prague, Czechoslovakia,
September 14--18, 1992.
\vspace{1pc}
]
..
\onecolumn

The \twocolumn command, called a “declaration” in Lamport’s book (p. 82), accepts an optional argument, enclosed in square brackets, which will be typeset in a single column. Closing the optional argument with the closing square bracket then yields two-column output until it is halted by the \onecolumn command.

Occasionally we run into trouble getting things to fit where we want them in this newsletter — no surprise here: it happens to all TEX users at some point or another. First came all the usual ploys: use a smaller font \small, and re-

duce the \baselineskip (down to 11pt). But the {enumerate} and {itemize}

environments left too much space in the left margin area; instead, I used list, which I found easier to work with since it requires the user to specify parameters, rather than dealing with redefining those already assigned to the other environments mentioned above. And it also gave a fairly dense layout, which was what I needed.

{\small \baselineskip=11pt
\begin{list}{$\triangleright$}{\itemsep=1pt\leftmargin=0pt
                               \itemindent=11pt}
   \item Alan Hoenig: When \TeX\ and \MF\ Work Together
         (invited talk) p.~1.
   \item John D.\ Hobby: Introduction to MetaPost (invited talk)
         p.~21.
\end{list}

This issue of making things fit is an “interesting” constraint, one which doesn’t arise all that often in one’s regular TEX activities. And it’s one which isn’t always handled, uh, “gracefully”, shall we say?

Peter Flynn forwarded a note he’d received regarding some spacing ugliness which had arisen in TTN 1,2 on p. 14, the para. beginning “Finally, I decided that … ”. This was particularly glaring since Peter’s own column a few pages before this had been going on about how poorly some other program did leading.

Peter very gallantly said it was probably so that I could get everything to fit, and I of course hastily agreed — “that’s what it must’ve been.”

Well … no. I went and looked at the file. This hadn’t been the result of willful squishing — I had made an error in using L^{A}TEX, one which I suspect is not unique to me:

   Finally, I decided that .... came up with the following:
   \vspace{-6pt}
   {\small
   \begin{verbatim}
   ..........

By not leaving a blank line between the end of the para. and the \vspace command, the leading normally associated with \small was affecting the text above it. To test this, I put a blank line above \vspace{-6pt}, and the paragraph’s normal leading was returned — with no change to the page break. So — user error. And while we do need to pack a lot of material into our 28 pages, it’s preferable that it be done on purpose and not by accident!

AD COMES HERE

(L^{A})TEX News

Volunteer work for the L^{A}TEX3 project

Frank Mittelbach, Chris Rowley, Michael J. Downes

This is a call for volunteers to help in the development of L^{A}TEX3. A number of these specialized tasks can be worked on concurrently with the development of the L^{A}TEX3 kernel; and some require special expertise not found among the core programming team. Initial research, analysis, and work on these tasks by volunteers can greatly speed up the process of integrating a number of desirable features into L^{A}TEX3. Many of these features can be extensively developed and tested under L^{A}TEX 2.09 even before the L^{A}TEX3 kernel is available.

Therefore we are publishing a list of tasks to the L^{A}TEX user community through various channels. The task list is available via anonymous ftp from the following sites: niord.shsu.edu, directory [fileserv.vol-task];

ftp.uni-stuttgart.de, directory soft/tex/vol-task. For access via mail

server, send mail to fileserv@shsu.bitnet, with no subject line and in the body

write: sendme vol-task, or to mail-server@rus.uni-stuttgart.de, with no
subject line and in the body write: send soft/tex/vol-task/vol-task.tex.

As well, it will appear in TUGboat 13, no. 4. If you are unable to retrieve a copy via electronic networks, please contact: L^{A}TEX3 Volunteer Task List Request, c/o TEX Users Group, P.O. Box 9506, Providence, RI 02940 USA.

The task list will be updated at regular intervals. Some of the main tasks include:

• Validating L^{A}TEX 2.09.

• Syntax questions: .sty metacomments for smart editors; syntax proposal for bibliographical commands; research on syntax for tables, for chemistry, and for commutative diagrams.

• Research tasks: experimenting with \emergencystretch; research on indexing commands; research footnote/endnotes conventions; syntax diagrams; BNF notation. Also: research on use of shorthand forms; research on figures and captions; research on the use of ^^conventions.

• Miscellaneous items: math font handling; converting numbers to textual form; rewrite of MakeIndex in WEB; write other auxiliary programs; bibliography style programming; bibliography style requirements; timing tests.

News from Around …

Gleanings from here and there

Harumi Fujiura wrote that the usual August meeting of the Japanese TEX user group didn’t take place — not enough papers.

The Polish user group, GUST, held a meeting at the beginning of this month, October; Hanna Kołodziejska says they’re looking into starting up an archive, Włodek Bzyl at the University of Gdańsk is working on the first issue of their newletter, and there may be a TEX meeting sometime next spring.

The UKTEXUG held its Annual General Meeting on 14 October at Aston. We should be getting a report for the next issue of TTN.

CyrTUG was unable to hold its first annual meeting last May due to internal upheavals in the country. Irina Makhovaya tells me that the meeting will be in a few days’ time, October 20–22; she’ll send us a short report.

The Dutch user group, NTG, is having its 10th meeting in Meppel (near Groningen) beginning at 14:00, with the theme “L^{A}TEX, L^{A}TEX3 and Font Selection”. Speakers will include Frank Mittelbach and Chris Rowley, Yannis Haralambous, and Johannes Braams. There are plans to visit a typographics museum in the morning. Contact Gerard van Nes at vannes@ecn.nl for details.

The Nordic group is planning a meeting also in November, date yet to be determined. Contact Roswitha Graham (see the Upcoming Events column for addresses).

Enn Saar of the Estonian user group wrote to say that they’re “still waiting” for ftp access to be set up. They’ve got a book on TEX at the publishers, and will send us information as soon as it’s available.

Reports on Meetings

TUG’92: 27–30 July Portland, Oregon

Christina Thiele

The 1992 TUG meeting in Portland, Oregon was, to my mind, a great success. Of the 141 pre-registered attendees, 107 were from the States, fully 21 people had come from Europe, 10 from Canada, and three from elsewhere (Israel, Japan, and Venezuela!). By the time the meeting had begun, there were some 150 in attendance.

The weather was unusually hot and dry for Portland at this time of year (an official drought had, in fact, been declared), but the newly renovated Benson Hotel was a wonderful pool of cool air, charming old-style atmosphere, and very good meeting rooms.

Oh, the conference? Well, since this was the first time since 1987 that I’d been able to sit back and and attend the papers, not having to worry about the program or the schedule or the unexpected, it was terrific. The first day, with papers and workshops on wrestling with PostScript and TEX, was of particular interest since I’ll be moving in that direction in the next half year — starting with Malcolm Clark’s keynote address, on through five papers, and ending with Anita Hoover’s workshop on PostScript, there’s plenty of material to read through and study in the Proceedings. And since I’ve mentioned the Proceedings, I think I’ll simply refer you to its pages, rather than provide a blow-by-blow account of each presentation. Other workshops covered such issues as archive management, standard coding in production, and modifying L^{A}TEX style files and manmac.sty.

To my mind, the highlight — both in terms of the presentation and the audience’s reaction — had to be the paper by Raman V.T.: “An Audio View of TEX Documents”. Raman is blind, and uses a voice synthesizer to translate TEX document structure into speech. His demonstrations left us, the audience, rather speechless — especially when he then showed how non-speech sounds (pauses, tone sounds of various pitch and length) could be used to represent such elements in mathematical expressions as operators, braces, and so on. Listening to his taped examples, I could only laugh at my own shortsightedness — why not use sounds for operators and other items which, when spoken out loud, must be referred to by their often unwieldy labels: “open parenthesis,” “multiplied by,” “squared,” and so on. It was like opening an entirely new section of the library (you see how one is tied to the written image of language?!) and discovering TEX all over again.

The afternoon was given over to a panel discussion on archives, the TUG business meeting, and then a special dinner held in Atwater’s Restaurant on the 41st floor of the U.S. Bancorp Tower — windows looking out in three directions. The view was quite breathtaking — Mount Hood to the east, and Mount Saint Helens’ shattered cone visible to the north — and as the evening became dark, the lights came on in the city below. But it wasn’t all splendour and awe — the Program Committee had found a way to keep us from reverting to shop talk: a magician! It was simple fun, but it was such good fun. Photos will surely circulate, but I won’t give anything away here!

For several years now, Malcolm Clark and Doug Henderson have struggled with the great uninformed masses who have not appreciated the finer points of bowling. It was tried at College Station … and again in Cork (yup, we actually went bowling in Cork!). But the Wednesday evening bowling tournament to raise funds for the L^{A}TEX3 project must now stand as the pinnacle, the culmination of their years in the wilderness … well, in the bowling lanes, at least.

By Thursday, many of us were almost (but not quite) talked and TEX’d out. Finally (some would say, mercifully) by around 3:30, it was over. People slowly collected their papers, their flyers, their books (should I mention that Portland boasts an enormous bookstore called Powell’s … I’m not kidding when I say we all left with much more than we’d arrived with!), their notebooks, their preprints; some left that evening; some, like myself, only left on Friday. It was a slow disappearing act.

We had had so much to listen to, so many people to talk to, Portland the city was rather lost in our schedules. But it was a beautiful city for a conference, and many people were using it as a starting point to their vacations.

I would like to congratulate the Program Committee of Doug Henderson, Helen Gibson, and Ron Whitney, with Mimi Jett as Program Coordinator, on having put together a very good conference, an enjoyable conference, a busy and stimulating conference. And so the challenge now goes to the next site — Aston University, Birmingham, UK — and to Peter Abbott. Shall we bowl again? … I’ve heard mutterings about darts … We have a year to wait. A year to plan. Only about half a year to write a paper! So go circle the last week of July 1993 on your calendars!

Bowling Tournament Boosts LATEX3 Fund

Doug Henderson Co-Organizer

On Wednesday evening, we held a L^{A}TEX3 Bowling Fundraiser at the Portland TUG meeting. Twelve four-person teams took to the lanes of glory to bowl for a most noble cause. Supporting further research for the L^{A}TEX3 system was our goal, but the rallying effort of the fun-loving conference attendees was amazing. We had a full one third of the conference down at the Hollywood Bowl to watch as people hurled that heavy ball (and sometimes themselves) down the lane to prove to the world that much fun can be had by all for a good cause.

It was a fun tournament. My fondest memory will be my worries about the scoring of twelve different teams and the “Tournament Chart” we were trying to construct before the event that evening. I scratched my head. No help. We abandoned our half-made chart in favor of winging it. When we arrived and talked with the nice folks behind the counter, they said “No problem. We have computerized scoring and can set you up with a tournament package for twelve teams.” Instant relief. Of course, it was too good to be true: we got fouled up in scoring who was where, and the computerized system rested while we applied the time-tested pencil and paper tallying techniques.

When all was said and done, the Blue Sky Research bowling team won the high-pin team total, but they were disqualified for having too professional a bowling shirt! The tournament winners (presented with real bowling trophies) were the team of: Chris Carruthers, Tom Thompson, Mary Louise Garcia, and Dave Steiner. The second-place team (presented with second place ribbons) included: Ron Whitney, Dan Olson, Brett Perkes, and Dave Marks. All those who participated were winners in my view, and we all had big smiles as we made the presentation/donation to Chris Rowley, the United Kingdom L^{A}TEX3 team member who was present. Our fundraiser brought in $740 for L^{A}TEX3. The following note, listing all the participants, was typeset (of course it was done in TEX) and presented along with the funds.

“We, the participants, would like to offer our support to the L^{A}TEX3 project. Please accept this contribution in the spirit intended. We had fun, we bowled until we were ready to drop, and there were no overfull boxes. Thank you for undertaking such an important project. Continued good luck and best wishes.”

Euro-TEX’92: 15–18 Sept. Prague, Czechoslovakia

Alan Hoenig

Member, TUG Board of Directors

Elsewhere in this issue, you will find a list of the presentations at the recent 7th EuroTEX Conference held in Prague from September 15 to 18. It is my pleasant task to here flesh out that list with some personal observations.

First of all, a hearty shame to those readers who were not there. The entire bill of fare for the meeting — including lodgings, meeting fee, pass for all Prague public transport during the meeting, and one or two other nice things — was about 300 DM (≈ $200 or ≈ £100). It was a particular shame that more people from the Americas were not present. Despite the slightly pricey airfare to Prague, surely the cost of the entire week was comparable or even less than that of a professional meeting in some distant part of the States.

But there is a double shame, for the meeting was run with a degree of efficiency and professional élan not always in evidence at meetings like this. The program organizers — Jiří Veselý, Jiří Zlatuška, and Karel Horák — crossed every t and háček’ed every ř, from the tours organized for the loved ones of the attendees to the meeting Proceedings, which were hardbound and waiting for us on the last day of the conference. Of course, the social events that attendees have come to love and anticipate helped fill the evenings.

I will not say much more about the actual program, for brief comments can only short change the presenters. I urge instead that you seek out the Proceedings. It has been recent custom to award a prize to the “best” presentation, where “best” is an idiosyncratic notion governed by someone in charge of the current conference. This prize has come to be known as the Cathy Booth Prize in honor of the late leader of the user community, and this year’s recipients were Bogusłav Jackowski and Marek Ryćko for their work in the development of Polish TEX.

It’s personally exciting to me to see the particular enthusiasm that Europeans and Eastern Europeans bring to TEX and . As usual, there was intense interest in the emerging DC font standards. Some of the work the Czechs are doing with is noteworthy, including Karel Horák’s use of it for mathematical figures and the -specific editor under development by a group of students under the direction of Dr. Mirko Křivánek.

It would not be right to conclude an account of a meeting in Prague without some mention of the golden city itself. This was my first time in Eastern Europe, and what a remarkable initiation! The central city is truly a story book set. Public spaces like Old Town Square and the magnificent Charles Bridge keep inviting the visitor back for yet another stroll. The Czechs were gracious and friendly, and it was easy to find someone who understood the English, pidgin German, or Russian that my wife or I were able to muster up. Public transport was clean and ultra-reliable, and the transport pass issued to all guests meant we could wander back and forth throughout the city at will, courtesy of the bus, tram, and subway systems. We felt safe strolling about at any time of day or night.

I know there are many other fine spots in Europe to serve as sites for upcoming EuroTEX conferences, but I look forward to the time when it’s Prague’s turn again.

TUG Board Activities

Summary of Board Meeting

The TUG Board of Directors met for two days before (25–26 July) and one afternoon (30 July) after the Annual Meeting of the TEX Users Group, in Portland, Oregon.^{1}

Following welcoming remarks by Malcolm Clark, TUG’s President, the preliminaries to the meeting were dealt with: reviewing and prioritising the agenda, adding new items, scheduling committee meetings, and accepting the minutes (with some corrections) from February’s board meeting.

24 TEX and TUG NEWS Vol. 1, No. 3, 1992

Between February’s board meeting and July, three motions had been voted upon by e-mail, and were ratified at the present board meeting: Jon Radel was re-appointed Public Domain PC Software Coordinator; a proposal by Peter Flynn for a lottery was defeated; and a bylaws change to TUG’s status (from 501(c)3 to 501(c)6) was accepted.

1Present: Peter Abbott, Barbara Beeton, Malcolm Clark (Chair, TUG President), Luzia Dietsche, Ken Dreyhaupt (Vice-President), Michael Ferguson, Peter Flynn (Secretary), Yannis Haralambous, Doug Henderson, Alan Hoenig (absent on Saturday), Anita Hoover, Mimi Jett, David Kellerman, Joachim Lammarsch, Jon Radel, Christina Thiele, Bill Woolf (Treasurer); Ron Whitney, TUG’s Business Manager, was also present. Absent: Bernard Gaulle, Roswitha Graham, Kees v.d. Laan, and Nico Poppelier.

Discussions on the election procedures for president included distribution of the nomination forms, an improved schedule for returning ballots, and an outline of a change to the bylaws to provide for a 2 1/2 year term (in order to bring the term of president into synch with those of the board).

The Bylaws Committee reported on revisions and improvements they had identified since being charged with this task at Aston. In connection with the previous item, one major change was to transfer all explicit references to elections (both for the board and the position of president) to a new Elections Procedures document. After discussion of these items, and suggestions on phrasing, the proposed changes were adopted. Minor changes in wording regarding candidates for election were also accepted unanimously.

On both Saturday and Sunday, the board spent much time in general discussions on various aspects of the operation and role of the TUG Office and of TUG. Issues ranged from what the main office priorities should be to the importance of the L^{A}TEX3 project on TEX’s continued existence. Generally, it seemed there was consensus on the priorities of membership services and volunteer coordination for the office. Mimi Jett had prepared a report which the Executive would be able to use in its search for a new Executive Director. A Search Committee, comprising the Executive and Mimi Jett was established. In all, it was a very useful exercise, providing background information to the newer board members and allowing all board members to express their perspectives.

Discussions on the budget included the final figures for 1991, which showed a loss of $24,454. Ron Whitney went over various points in the budget for 1992 (projected deficit of $3,100), and answered questions the board had on several items. A 1993 budget had also been prepared in outline, based on the 1992 information. Additional budget items were pencilled in as the board made various decisions with financial implications: membership dues for 1993 were to remain at 1992 levels; student membership would be set at 50% of regular membership, beginning in 1993; allocations for various committees were also approved (e.g., Knuth Scholarship). Costs related to the arrival of a new Executive Director would have to be added when they became known.

Reports from various committees were presented, some on their work in progress, others with final reports requiring action. The Long Range Planning Committee reported that their work was progressing slowly, and that a draft document would be ready for the fall. The Committee on Special Vice-Presidents (Luzia Dietsche, chair) presented a number of proposals which included both a statement of intent and some suggested bylaws amendments. The statement of intent was passed, and the proposed bylaws amendments were referred to the Bylaws Committee for implementation.^{2} The Conference Planning Committee (Christina Thiele, chair) had previously circulated their report, proposing Aston University, at Birmingham, UK, be accepted as the site for the 1993 TUG Annual Meeting; Peter Abbott agreed to undertake the arrangements. With regard to the future, several sites had already expressed interest in the 1994 meeting, and it was hoped that new procedures on longer-term site selection would be implemented soon. Mimi Jett (Program Committee) provided a brief run-down of the Portland meeting and the events planned. Barbara Beeton (Knuth Scholarship liaison) announced the winner of the 1993 scholarship (Jenny Smith of Chichester, UK). All committees were reviewed and repopulated where necessary.^{3} New committees established included the already-mentioned Search Committee, and a Promotions Committee.

2The statement of intent reads, in part: “The term “Special Vice-Presidents” shall be replaced by “Special Director”. A Special Director shall have all the rights and responsibilities of a Director except that Special Directors will not be counted in determining a quorum. A Special Director shall be invited by the TUG Board as the representative, usually the President, of a Local TUG.

An area where there was increasing need to respond by TUG involved the support which could be provided new user groups, particularly in Eastern Europe. Yannis (chair of the committee) presented a number of proposals, in particular the reprinting of a facsimile of TUGboat, in order to make it available more cheaply: currency problems are a major hindrance to taking out a membership in TUG. Yannis was going to find out more details from people involved in the various groups who would be attending the Prague EuroTEX’92 meeting.

Michael Ferguson presented an overview of the Technical Council, and Alan Hoenig reported on his work as liaison with the L^{A}TEX3 project. Proposals for bylaws amendments were not accepted, on procedural grounds, and were referred to the Bylaws Committee. A proposal from Sebastian Rahtz regarding archives was circulated to the board, and passed along to the Technical Council for further discussion. Also referred to the Technical Council was recent correspondence from Prof. Knuth, regarding the registration of several icons for TEX and related programs with SUN.

Other issues addressed by the board included board e-mail communications, the “Aston Shop” (TUG articles for sale via Peter Abbott), site coordinators (perhaps become part of the Technical Council), the resource directory, departing gifts to members of the previous board (those who did not run for election under the new rules), TUG’s tax exempt status, and TUG investments. There were also some discussions on various aspects of the Portland meeting, which had been applauded most favourably by members of the board.

TUG Presidential Election ^{Malcolm Clark}

TUG President

No nominations for TUG President were received as of the deadline, September 28, 1992.

The TUG Bylaws (Article VII, Section 10) state that “When the office of President becomes vacant, the Vice President will become President for the remainder of the President’s term and will then, as President, appoint a member to serve as Vice President.” Thus, if nothing else happens, the position of President will fall vacant on January 1, 1993, and the person holding the office of Vice President on that date will become President.

It is the concern of all that the person to become President be someone willing and able to spend the time and energy necessary to perform the duties of the office. Several natural possibilities — including the current Vice President — have indicated that finding the time would be difficult. The Board is reviewing the options available to TUG and will report to the membership in the next issue of TTN (which is scheduled for mailing in early December).

3See latest issue of TUGboat for listing.

TEX Users Group 1993 Course Schedule

Beginning/Intermediate TEX* Intensive Course in L^{A}TEX*
Boston  April 19–23  San Francisco March 1–5
San Diego  June 14–19  Boston  April 26–30
Boston  August 9–13  Ottawa  August 23–27
Chicago  October 18–22  Boston  October 25–29
Modifying L^{A}TEX Style Files* Adv’d TEX and Macro Writing*
San Diego  June 7–11  San Diego  June 21–25
Boston  November 1–5
TEX for Publishers  Practical SGML and TEX*
Boston  February 23  San Francisco March 8–9
Washington, DC April 14  Boston  November 8–9
New York City June 9
Orlando  September 22

Book and Document Design with TEX

Boston September 23–24

*Lab classes — computers will be provided for all students.

• TUG courses are small, with 8–15 students in most classes and are held at major hotels.

• The dates and locations above are tentative — watch for the final schedule in the mail this fall.

• For more information, contact the TEX Users Group at (401)–751–7760.

On-Site Courses in TEX and L^{A}TEX from TEX Users Group

• Courses in TEX, L^{A}TEX, SGML and TEX, PostScript, or TEX for Publishers tailored to the needs of your group

• Courses at every level from beginning to advanced

• Five full days of instruction at your site

• One-week course fee includes all instructor fees and expenses plus textbooks and other materials for up to 15 students

• If a properly equipped training facility is not available, TUG will arrange computer rentals and supply TEX or L^{A}TEX software

Upcoming Events

November Nordic Group: Annual General Meeting. Copenhagen, Denmark. [Date to be announced] Roswitha Graham

                           roswitha@admin.kth.se

Peter Busk Laursen

                           unipbl@uts.uni-c.dk

19 November

NTG:

“L^{A}TEX, L^{A}TEX3 and Font Selection.” Meppel (near Groningen), the Netherlands.

Gerard van Nes

                           vannes@ecn.nl

23–26 Nov.

CNRS:

“Traitement de texte scientifique sur Macintosh.” Ivry, France.

K.B. Su, Chercheur CNRS, Laboratoire d’electrochimie, Université Paris 7 Tél: 49.60.40.29, 49.60.40.30

{24–27 Feb. \atop 1993}

CONCEPPTS’93:

The Prepublishing Conference (an international conference on computers and electronic publishing and printing technologies). Orange Co. Convention Centre, Orlando, Florida.

Phone: 703-264-7200 FAX: 703-620-9187.

9–12 March DANTE’93 and General Meeting, Chemnitz (near Dresden).

Dr. Wolfgang Riedel

                           wolfgang.riedel@hrz.tu-
                           chemnitz.de

26–29 July

TUG’93:

“World Wide Window on TEX” Aston University, Birmingham, UK.

Chris Rowley, ca rowley

                           @vax.acs.open.ac.uk

Malcolm Clark

                           malcolmc@wmin.ac.uk

Note: Also consult the “Calendar” in the next issue of TUGboat for more dates.

EuroTEX’92: Contents from the Proceedings

The following is excerpted from the Proceedings of the 7th European TEX Conference, held in Prague, Czechoslovakia, September 14–18, 1992.

▷ Alan Hoenig: When TEX and - Work Together (invited talk) p. 1. ▷ John D. Hobby: Introduction to Meta-Post (invited talk) p. 21. ▷ Andrei B. Khodulev and Irina A. Makhovaya: On TEX Experience in Mir Publishers p. 37. ▷ Laurent Siebenmann: The Lion and the Mouse p. 43. ▷ Petr Sojka, Rudolf Červenka, and Martin Svoboda: TEX for Database Publishing p. 53. ▷ Ondřej Vácha: TEX for Typesetting in a Publishing House p. 59. ▷ Michel Lavaud: A Solution to Help Ensure the Future of TEX: Make its Use Easier on Cheap Machines p. 66. ▷ Frank Mittelbach and Chris Rowley: L^{A}TEX3: Structure and Design (invited talk) p. 69. ▷ Anita Z. Hoover: The Key to Successful Support: Knowing Your TEX and L^{A}TEX Users (invited talk) p. 71. ▷ Theo Jurriens: TEX for Everybody? p. 86. ▷ Daniel Flipo and Laurent Siebenmann: Hyphenation in Presence of Accents and Diacritics: An Easy and Low-Cost Solution p. 87. ▷ Laurent Siebenmann: CaesarCM—A Gentle Road to Perfect Hyphenation in Modest TEX Environments p. 97. ▷ Laurent Siebenmann: La mise en application définitive de la norme de Cork p. 107.

▷ Bogusłav Jackowski and Marek Ryćko: Polishing TEX: From Ready to Use to Handy in Use p. 119. ▷ Rama Porrat: Developments in Hebrew TEX p. 135. ▷ Michael Vinogradov: Russian TEX: New Eight-Bit Fonts and IBM PC Equipment p. 149. ▷ Klaus Lagally: ArabTEX—Typesetting Arabic with Vowels and Ligatures p. 153. ▷ Erik-Jan Vens: Incorporating PostScript Fonts in TEX p. 173. ▷ Kristoffer H. Rose: How to Typeset Pretty Diagram Arrows with TEX: Design Decisions Used in XY-pic p. 183. ▷ Kees van der Laan: Table Diversions p. 211. ▷ Jörg Knappen: Changing the Appearance of Math p. 212. ▷ Kees van der Laan: Typesetting Crosswords via TEX p. 217. ▷ Kees van der Laan: fifo and lifo Incognito p. 225. ▷ Philip Taylor: The Future of TEX p. 235. ▷ Frank Mittelbach and Chris Rowley: The Future of High-Quality Typesetting: Structure and Design (invited talk) p. 255. ▷ Daniel Taupin: MusicTEX: Using TEX to Write Polyphonic or Instrumental Music (invited talk) p. 257. ▷ Kristoffer H. Rose: Typesetting Diagrams with XY-pic: User’s Manual p. 273. ▷ Yannis Haralambous: Towards the Revival of Traditional Arabic Typography … through TEX (invited talk) p. 293.

TEX and TUG NEWS Table of Contents

Editorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 The Donald E. Knuth Scholarship

Barbara Beeton, Nico Poppelier . . . . . . . . . . . . . . . . . 3 L^{A}TEX3 Update

Frank Mittelbach, Chris Rowley, and Alan Hoenig . . . . . . . 4 The Technical Council

Michael Ferguson . . . . . . . . . . . . . . . . . . . . . . . . . 5 New Publications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Typographer’s Inn

Peter Flynn . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 “Hey — it works!” . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Marking lines in the margin Peter Schmitt . . . . . . . . . . . 12 When there’s only one of you and

hundreds of them Mimi Burbank . . . . . . . . . . . . . 14 Some TTN L^{A}TEX tricks Christina Thiele . . . . . . . . . . . . 15

(L^{A})TEX News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Volunteer work for the L^{A}TEX3 project

Frank Mittelbach, Chris Rowley, Michael J. Downes … 18

News from Around . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Gleanings from here and there . . . . . . . . . . . . . . . . . . . 19 Reports on Meetings . . . . . . . . . . . . . . . . . . . . . . . . . . . .

TUG’92 (Portland) Christina Thiele . . . . . . . . . . . . . . . 19 Bowling Tournament Boosts L^{A}TEX3 Fund Doug Henderson . . 21 EuroTEX’92 (Prague) Alan Hoenig . . . . . . . . . . . . . . . . 22

TUG Board Activities . . . . . . . . . . . . . . . . . . . . . . . . . . . Summary of Board Meeting . . . . . . . . . . . . . . . . . . . . . 23 TUG Presidential Election . . . . . . . . . . . . . . . . . . . . . . 25

TEX Users Group 1993 Course Schedule . . . . . . . . . . . . . . . . . 26 Upcoming Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 EuroTEX’92: Contents from the Proceedings . . . . . . . . . . . . . . 28

Volume 1, No. 3, 1992

The latest TEXniques volumes from TUG

A Gentle Introduction to TEX by Michael Doob. TEXniques 13. Providence, Rhode Island: TEX Users Group, 1990. vi + 88pp. (softcover), US$15. $13 for TUG members.

• The book provides a first look at TEX for those who have little experience with the program. It begins with the basics (characters, fonts, control sequences), and then moves to parameter settings for control of line, paragraph, and page layouts. Groups, basic math constructs and tables are treated, along with macro definitions and practice with error messages.

METAFONTware by Donald E. Knuth, Tomas G. Rokicki, and Arthur L. Samuel. TEXniques 14. Providence, Rhode Island: TEX Users Group, 1990. iii + 101pp. (softcover), US$20. $18 for TUG members.

• This report comprises the WEAVEd output of the four programs GFType, GFtoPK, GFtoDVI, and MFT. The listed programs are current as of February 1990 and incorporate changes corresponding to those in version 1.8, including support of 8-bit input and extended ligatures.

A Permuted Index for TEX and LATEX by Bill Cheswick. TEXniques 15. Providence, Rhode Island: TEX Users Group, 1991. 111pp. (softcover), US$15. $13 for TUG members.

• This index is designed to help TEX and L^{A}TEX users find the right command among the one thousand-odd commands documented. A description of the method used to construct this index is given in TUGboat 10, no. 4, pp. 707–713.

All issues of TEXniques are available from the TUG office (see list published on yellow order form at the back of TUGboat 13, no. 1).

TEX Users Group P.O. Box 9506 Providence, RI 02940 USA

                    tug@math.ams.org